In a study of pain relievers, 100 people were given product A, and 67 experienced relief. In the same study, 25 people were given product B, and 20 experienced relief. Fill in the blanks of the statement below to make the statement the most reasonable possible. Product ?AB performed better in the study because only % failed to get relief with this product, whereas % failed to get relief with Product ?AB
Answer #1

With product A, 100 - 67 = 33 failed to get relief

Percentage who didn't get relief = (33/100) * 100 = 33%

With product B, 25 - 20 = 5 failed to get relief

Percentage who didn't get relief = (5/25) * 100 = 20%

Product B performed better in the study because only 20% failed to get relief with this product, whereas 33% failed to get relief with Product A.